### v4

# Hurst index of gamma-ray burst light curves and its statistical study

Abstract
Gamma-ray bursts (GRBs) rank among the most powerful astrophysical phenomena, characterized by complex and highly variable prompt emission light curves that reflect the dynamics of their central engines. In this work, we analyze a sample of 163 long-duration GRBs detected by the Burst and Transient Source Experiment (BATSE), applying detrended fluctuation analysis (DFA) to derive the Hurst index as a quantitative descriptor of temporal correlations in the light curves. We further explore statistical correlations between the Hurst index and 12 other observational parameters through regression and correlation analyses. Our results reveal anti-correlations between the Hurst index and the burst durations ($T_{50}$, $T_{90}$), and moderate positive correlations with peak photon flux proxies (Ppk1–Ppk3). By contrast, the standard spectral parameters (including the low-energy index $\alpha$) show no evidence for a linear dependence on the Hurst index in our sample. We do not find a clear monotonic weakening of the correlation strength from 64 ms to 1024 ms peak-flux measures; rather, the correlation coefficients for Ppk1–Ppk3 are comparable within uncertainties. The results offer new perspectives on the temporal structure of the GRB emission and its potential link to the underlying physical mechanisms driving these bursts.
keywords:
Gamma-ray bursts (629) , Time domain astronomy (2109)
journal: Journal of High Energy Astrophysics
[…]
[label1]organization=Department of Astronomy, School of Physics, Huazhong University of Science and Technology,city=Wuhan, postcode=430074, country=China
[…]
[label2]organization=School of Mathematics and Physics, Qingdao University of Science and Technology,city=Qingdao, postcode=266061, country=China
1 Introduction
Gamma-ray bursts (GRBs) are the most powerful and luminous transient events in the universe, believed to originate from relativistic jets produced by catastrophic astrophysical processes. Observationally, GRBs are broadly classified into two categories—long and short—based on a characteristic duration threshold of approximately two seconds(Kouveliotou et al., 1993). Long GRBs ($T_{90}>2\,\mathrm{s}$) are generally associated with the core collapse of massive, rapidly rotating stars(Woosley, 1993; Paczyński, 1998) and are frequently accompanied by broad-lined Type Ic supernovae(Woosley and Bloom, 2006). In contrast, short GRBs ($T_{90}<2\,\mathrm{s}$) are commonly linked to the mergers of compact binary systems(Li and Paczyński, 1998), such as neutron star–neutron star(Abbott et al., 2017; Goldstein et al., 2017; Paczynski, 1986) or neutron star–black hole systems(Meszaros and Rees, 1992; Paczynski, 1991).
The light curves (LCs) of the GRBs serve as a vital diagnostic of the underlying physical processes governing these extreme astrophysical explosions and the conditions in the vicinity of their central engines. Characterized by highly variable, non-thermal emission that spans timescales from milliseconds to several minutes (MacLachlan et al., 2012), GRB light curves display a remarkable range of temporal structures, including single-peaked components, multi-episodic pulses, and intricate patterns of variability across different energy bands. This morphological diversity indicates the complex nature of relativistic outflows and energy dissipation mechanisms, which may involve internal shocks, magnetic reconnection, or cascades driven by turbulence within the jet (Zhang and Yan, 2011).
Over the years, numerous indicators have been developed to quantify the temporal variability of GRBs. These include pulse decomposition analysis (Norris et al., 1996), variability indices (Fenimore and Ramirez-Ruiz, 2000b), Fourier power density spectra (Beloborodov et al., 2000), autocorrelation functions (Borgonovo, 2004), and minimum variability time scales (MacLachlan et al., 2012). Crucially, some studies have found significant correlations between variability indicators and other properties such as luminosity (Reichart et al., 2001; Guidorzi et al., 2005), suggesting that the light curve structure may provide key diagnostics of jet composition and energy dissipation.
Originally introduced by Peng et al. (1994, 1995), detrended fluctuation analysis (DFA) is a fundamental method for examining the scalar properties in a variety of time series datasets. Later, Kantelhardt et al. (2002) expanded DFA to analyze multifractal processes, resulting in the multifractal detrended fluctuation analysis (MFDFA) technique. A generalized Hurst index was defined to quantify the long-range dependence of a time series, initially introduced by Hurst (1951) in the study of the long-term storage capacity of reservoirs. This method effectively addresses correlation issues in time-series data and is applicable to both discrete and continuous stochastic processes. It has been used successfully in a wide range of applications spanning several fields, such as music (Jafari et al., 2007), heartbeat dynamics (Ivanov et al., 1999), electroencephalogram sleep data (Pavlov et al., 2020b, 2021), arterial pressure (Pavlov et al., 2020a), cosmic microwave radiation (Movahed et al., 2011, 2013), atmospheric turbulence effects on stellar images (Zunino et al., 2014), sunspot fluctuations (Sadegh Movahed et al., 2006; Hu et al., 2009), solar flares (Lee et al., 2020), gravitational wave detection (Eghdami et al., 2018), quasiperiodic oscillation searching (Tarnopolski and Marchenko, 2021), blazars (Tarnopolski et al., 2020), and fast radio bursts (Wang et al., 2023). In this context, DFA and its extension to MFDFA offer powerful tools to characterize temporal correlations and long-range memory in GRB light curves. Unlike traditional variability measures, DFA quantifies the scaling behavior and persistence of non‑stationary time series, making it particularly suited to GRB data (Peng et al., 1994; Kantelhardt et al., 2002; Zunino et al., 2014).
Since its launch in 1991 onboard NASA’s Compton Gamma Ray Observatory, the Burst and Transient Source Experiment (BATSE) has been instrumental in shaping our current understanding of GRBs (Band et al., 1993). One of BATSE’s most impactful achievements is its systematic collection of high-resolution light curves in four well-defined energy bands: 20-50 keV, 50-100 keV, 100-300 keV, and above 300 keV, which has facilitated detailed multiband temporal and spectral investigations of GRB prompt emission. Using this extensive data set, Hakkila (2021) found that most GRB pulses can be well characterized by a smooth, single-peaked emission profile accompanied by a temporally symmetric residual component, shedding new light on the structure and physical origin of GRB emission mechanisms. We also chose the BATSE data for our analysis because of its high sensitivity.
In this paper, we analyze the detrended fluctuations in the light curves of 163 long GRBs detected by BATSE, derive a characteristic index (the Hurst index) from each GRB light curve treated as a time series, and perform regression and correlation analyses between this index and 12 different physical parameters that have been observed for these long GRBs. The paper is arranged as follows. In Section 2, we present our data sample and describe the 12 physical parameters for which we performed the correlation analysis. The detrended fluctuation analysis model, along with its parameter selection and statistical analysis, is presented in Section 3. The results and conclusions are reported in Sections 4 and 5, respectively.
2 Data Analysis
2.1 Sample Selection
Our analysis is based on a carefully selected sample of 163 long-duration GRBs detected by BATSE, corresponding to trigger numbers ranging from 107 to 1997111https://heasarc.gsfc.nasa.gov/FTP/compton/data/batse/ascii_data/64ms/, which is the intersection of the BATSE Gamma Ray Burst Lightcurve Image Archive [^2] https://gammaray.nsstc.nasa.gov/batse/grb/lightcurve/ and the CGRO/BATSE 4B Catalog [^3] https://heasarc.gsfc.nasa.gov/W3Browse/cgro/batse4b.html. The sample selection criteria also included: complete determinations of the spectral parameters from band function fittings and availability of all 12 other physical parameters under investigation from Wang et al. (2020). The 12 selected parameters are shown in Table 1. The DISCSC light curves are originally provided at 64 ms resolution. In this work, we rebin the DISCSC data to a uniform time resolution of 1024 ms by summing 16 consecutive 64 ms bins, so that all bursts are analyzed on the same temporal grid. This uniform binning is adopted because DFA results are sensitive to the sampling resolution; fixing the bin width across the sample avoids introducing an additional burst-dependent scale that could bias the distribution of Hurst indices and compromise the comparability required for our statistical study.
This choice also imposes a practical requirement on the minimum number of data points available for each burst. Starting from an initially larger candidate set, we excluded 36 GRBs with $T_{90}<8.6$ s, for which the rebinned light curves would contain too few 1024 ms bins within the prompt-emission window to yield a meaningful DFA scaling fit. After this filtering, 163 long GRBs remain and constitute our final sample. We note that relatively short bursts (e.g., those with $T_{90}\lesssim 30$ s) are still included as long as they satisfy the above minimum-length requirement under the uniform 1024 ms binning.
The time interval for each burst is selected from the trigger time up to $T_{90}$. This standardization provides a uniform and reproducible analysis window, guaranteeing that the scaling properties derived from detrended fluctuation analysis are representative of the burst’s intrinsic variability during its most active phase and enabling a meaningful comparison of Hurst indices for robust statistical evaluation. Fig. 1 presents the light curve of GRB 920110A as an example, one of the events included in our sample, over the time span corresponding to its $T_{90}$ duration. A (LABEL:tab:list_HI) lists the Hurst indices for all GRBs in our sample; the corresponding physical parameters used in the correlation analysis are taken from Wang et al. (2020). We emphasize that in this work $T_{90}$ is used as a practical duration parameter to define an analysis window that covers the vast majority of the prompt-emission variability, rather than as a strict physical gate that begins at the canonical $t_{\rm start}$ of the $T_{90}$ definition. In many BATSE long GRBs, the light curve shows a clear rise and non-negligible variability immediately after trigger ($t\approx 0$), and excluding the early segment prior to the onset of the cataloged $T_{90}$ interval would remove part of the burst structure that can contribute to the scaling behavior measured by DFA. Therefore, we adopt the interval $t\in[0,\,T_{90}]$ for all bursts to ensure methodological uniformity across the sample and to avoid selectively discarding early-time variability. The purpose of our study is to obtain a standard set of DFA-based scaling measures for statistical comparison across the sample. In Fig. 1, the dashed line marks the endpoint of the adopted analysis window at $t=T_{90}$ (measured from the trigger). The caption explicitly clarifies that $T_{90}$ is a duration parameter, and that the vertical marker indicates the end of the $0,\,T_{90}]$ window rather than the $T_{5},\,T_{95}]$ window. For five triggers (No. 107, 110, 114, 211, 351) whose publicly available DISCSC records terminate before $t=T_{90}$, we perform DFA on the actually available window $0,\,t_{\rm end}^{\rm data}]$ and treat the cataloged $T_{90}$ as a reference duration parameter.
Figure 1: Light curve of GRB 920110A as an example, the energy channel is 1-4 (>20>20 keV).
2.2 Physical Parameters
We analyze 12 key physical parameters that characterize different aspects of GRB emission. These parameters can be broadly categorized into three groups: (1) spectral parameters describing the energy distribution of photons, (2) temporal parameters characterizing the burst duration and variability, and (3) flux-related parameters measuring the burst intensity. Table 1 summarizes these parameters with their units and physical interpretations. The spectral parameters $\alpha$ and $\beta$ represent the low- and high-energy photon indices from the Band function fits, while $E_{peak}$ corresponds to the peak energy in the $\nu F_{\nu}$ spectrum. The temporal parameters $T_{50}$ and $T_{90}$ describe the durations of the bursts that contain total counts of 50% and 90%, respectively, and V quantifies the degree of fluctuations in the light curve. The flux-related parameters include fluence ($F_{g}$), peak flux ($F_{pk}$), and photon flux ($P_{pk}$) measured in different time bins. The hardness ratio (HR) provides information about the spectral hardness of the burst. All data were taken from Wang et al. (2020), while they were gathered from the references therein.
Parameter Units Description
$\alpha$ Opposite value of the low-energy spectrum index of the Band model
$\beta$ Opposite value of the high-energy spectrum index of the Band model
$E_{peak}$ keV Spectral peak energy of the Band model
$F_{g}$ $10^{-6}$ erg cm-2 Fluence in the 20-2000 keV energy band
$F_{pk}$ $10^{-6}$ erg cm-2 s-1 Peak flux with 1 s time bin in the observer’s frame 1-10000 keV energy band
HR Hardness ratio between 100-2000 keV and 20-100 keV
$P_{pk1}$ ph cm-2 s-1 Peak photon flux in the 64 ms time bin of 10-1000 keV
$P_{pk2}$ ph cm-2 s-1 Peak photon flux in the 256 ms time bin of 10-1000 keV
$P_{pk3}$ ph cm-2 s-1 Peak photon flux in the 1024 ms time bin of 10-1000 keV
$T_{50}$ s Burst duration between 25% and 75% of the cumulative counts (following Wang et al. (2020))
$T_{90}$ s Burst duration between 5% and 95% of the cumulative counts (following Wang et al. (2020))
V (variability) Light curve variability index based on Fenimore and Ramirez-Ruiz (2000a),
which also corresponds to variability1 in Wang et al. (2020)
Table 1: Description of the 12 physical parameters analyzed in this study. The parameters are grouped into spectral ($\alpha$, $\beta$, $E_{peak}$, HR), temporal ($T_{50}$, $T_{90}$, V), and flux-related ($F_{g}$, $F_{pk}$, $P_{pk1-3}$) categories. More detailed description can be found in Wang et al. (2020).
3 Methods
3.1 Detrended Fluctuation Analysis (DFA)
Detrended fluctuation analysis (DFA) was originally proposed by Peng et al. (1994) and later generalized to multifractal detrended fluctuation analysis (MF-DFA) by Kantelhardt et al. (2002). In this work, we focus on the monofractal case (DFA) by fixing $q=2$ in the MF-DFA formalism.
Given a GRB light curve represented by a discrete time series $x(i)$ of length $N$, DFA proceeds as follows:
(1) Profile construction. We construct the cumulative (integrated) profile
$y(l)=\sum_{i=1}^{l}\left[x(i)-\langle x\rangle\right],$ (1)
where $\langle x\rangle$ is the mean of the time series. Here $x(i)$ is the binned photon counts (counts per bin) in the $i$-th time bin of width $\Delta t=1.024~\mathrm{s}$ (after summing the four DISCSC channels).
(2) Segmentation and detrending. For a chosen scale (window size) $s$, the profile $y(l)$ is divided into $N_{s}=\lfloor N/s\rfloor$ non-overlapping segments of length $s$. To reduce boundary effects and to use the entire time series, we perform the same segmentation starting from the end of the profile, yielding a total of $2N_{s}$ segments (Kantelhardt et al., 2002). In each segment $\nu$, we fit an $n$-th order polynomial trend $y_{\nu}^{(n)}(l)$ (in this work $n=1$) via least squares and subtract it from the profile.
(3) Local variance. For each segment $\nu$, the detrended variance is computed as
$F^{2}(\nu,s)=\frac{1}{s}\sum_{k=1}^{s}\left[y_{\nu}(l)-y_{\nu}^{(n)}(l)\right]^{2},$ (2)
where $y_{\nu}(l)$ denotes the profile values restricted to segment $\nu$.
(4) Fluctuation function. The fluctuation function at scale $s$ is obtained by averaging over all $2N_{s}$ segments:
$F(s)=\left[\frac{1}{2N_{s}}\sum_{\nu=1}^{2N_{s}}F^{2}(\nu,s)\right]^{1/2}.$ (3)
(5) Scaling and slope estimation. Repeating the above steps for a set of scales $\{s\}$, DFA predicts a power-law scaling
$F(s)\propto s^{k},$ (4)
where $k$ is the fitted slope in the $\ln F(s)$$\ln s$ plane. We estimate $k$ using ordinary least squares (OLS) regression over all available scales used in the computation (see Section 3.2). To avoid confusion with the low-energy spectral index $\alpha$ of the Band function, we denote the DFA scaling slope by $k$ throughout this paper.
Following the convention implemented in the MFDFA package (Rydin Gorjão et al., 2022), we report the Hurst index for the light-curve series as
H=k1.H=k-1. (5)
Here $H$ is an empirical DFA scaling index defined by Eq. (5). For finite-length and strongly nonstationary burst light curves, the fitted scaling exponent may fall outside the canonical range expected for idealized fGn/fBm, and such cases should be interpreted with caution.
3.2 Implementation and analysis configuration
Time resolution and rebinning
The original BATSE light-curve data used in this work have a native time bin width of 64 ms. Prior to the DFA computation, we rebin each burst to a uniform time resolution of 1024 ms (1.024 s) by summing consecutive 16 bins. For a discrete count series $x_{64}(i)$ sampled at 64 ms, the rebinned series is
$x_{1024}(j)=\sum_{m=1}^{16}x_{64}\left(16(j-1)+m\right)\quad j=1,2,\dots,N_{1024}.$ (6)
All Hurst indices reported in this paper are derived from the rebinned 1024 ms light curves; the 64 ms data are used only as the input for rebinning. We note that BATSE monitors triggering on 64 ms, 256 ms, and 1024 ms timescales; this triggering timescale does not imply that our DFA is performed on 64 ms bins.
Time window definition
[…]
Scale selection and detrending order
We use logarithmically spaced scales from $s_{\min}=3$ bins to $s_{\max}=\lfloor N/10\rfloor$ bins, with 200 candidate scale points before integer rounding and deduplication. With $\Delta t=1.024$ s, the minimum scale corresponds to $s_{\min}\Delta t=3.072$ s. We adopt first-order polynomial detrending in each segment. For short rebinned series where the nominal upper scale $s_{\max}=\lfloor N/10\rfloor$ falls below $s_{\min}=3$ bins, we retain the burst by fitting over the available integer scales after rounding and deduplication.
Hurst index and uncertainty
For each GRB light curve, we evaluate the fluctuation function $F(s)$ on a set of logarithmically spaced window sizes $s$ and determine the scaling exponent by fitting the linear model
$\ln F(s)=k\,\ln s+b$ (7)
using an unweighted ordinary least-squares (OLS) regression over all adopted scales. We then report the Hurst index as H=k1H=k-1.
When the regression is well defined, the statistical uncertainty of the fitted slope $k$ is quantified by its OLS standard error. Specifically, let $n$ be the number of scale points used in the fit, $y_{i}=\ln F(s_{i})$, and $\hat{y}_{i}$ the fitted values. The residual sum of squares is
$\mathrm{RSS}=\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2},$ (8)
and the mean squared error (i.e., the unbiased estimator of the residual variance) is
$\mathrm{MSE}=\frac{\mathrm{RSS}}{n-2},$ (9)
where n2n-2 is the number of degrees of freedom for a linear fit (slope and intercept). With the design matrix $X=[\ln s_{i},\,1]$, the standard error of the slope is
$\mathrm{SE}(k)=\sqrt{\mathrm{MSE}\,\left[(X^{\mathsf{T}}X)^{-1}\right]_{11}}.$ (10)
Because $H$ differs from $k$ by a constant offset, the propagated uncertainty is simply $\mathrm{SE}(H)=\mathrm{SE}(k)$.
In a small number of cases (typically very short time series), the regression-based uncertainty may be ill-conditioned or numerically unstable. For these bursts, we estimate the uncertainty from repeated resampling of the time series (bootstrap or subsampling, depending on data length) and record the adopted uncertainty-estimation method for each GRB.
Worked example
To make the derivation of $H$ explicit, Figure 2 shows the $\ln F(s)$$\ln s$ scaling relation and the OLS fit for the GRB in Figure 1 (GRB 920110A). The corresponding $(s,F(s))$ data used for the fit are exported from our pipeline to enable full reproducibility.
Figure 2: $\ln F(s)-\ln s$ Scaling Relation and the OLS fit for GRB 920110A
4 Results
Based on the Hurst indices ($H$) measured for 163 BATSE GRBs (Section 3.2), we examine their statistical association with 12 fundamental GRB observables compiled by Wang et al. (2020). Following the approach of Wang et al. (2020) and the classical definition of the Pearson correlation coefficient (Pearson, 1895), we use the Pearson $r$ to quantify the strength of each HHxx relation. Beyond the point estimate of $r$, we (i) propagate measurement uncertainties to obtain an uncertainty interval for $r$, (ii) assess the adequacy of a linear description, and (iii) evaluate whether a linear trend is preferred over a null (no-dependence) model.
The uncertainty in $r$ is propagated from measurement errors via Monte Carlo (MC) sampling. For each observable $x$ and the Hurst index $H$, we generate $N_{\rm MC}$ synthetic realizations of the dataset by drawing, for each burst, $x_{i}^{\prime}\sim\mathcal{N}(x_{i},\sigma_{x,i})$ and $H_{i}^{\prime}\sim\mathcal{N}(H_{i},\sigma_{H,i})$, and compute the corresponding Pearson coefficient $r^{\prime}$ for each realization. We then summarize the resulting $r^{\prime}$ distribution by reporting its 16th and 84th percentiles as the central 68% interval, which we adopt as the MC-based uncertainty range of $r$.
To test whether a linear relation provides an adequate description of the data, we perform a linear regression that accounts for uncertainties in both coordinates and quantifies the goodness of fit using the reduced chi-square, $\chi_{\nu}^{2}$. Finally, to compare a linear trend against a null model in which $H$ is independent of $x$, we compute the Akaike and Bayesian information criteria (AIC and BIC) for both models and report the differences
$\Delta\mathrm{AIC}\equiv\mathrm{AIC}_{\rm null}-\mathrm{AIC}_{\rm lin},\qquad\Delta\mathrm{BIC}\equiv\mathrm{BIC}_{\rm null}-\mathrm{BIC}_{\rm lin},$ (11)
so that $\Delta\mathrm{AIC}>0$ and $\Delta\mathrm{BIC}>0$ indicate a preference for the linear model over the null model after accounting for model complexity. The results are presented below.
For the spectral/flux-related parameters $\alpha$, $\beta$, $\log E_{\rm peak}$, $\log F_{g}$, $\log F_{pk}$, and $HR$, the corresponding scatter plots are shown in Figs. 38. The corresponding Pearson’s $r$ values are close to zero and the information criteria do not favor adding a slope (Table 2), indicating that the data provide no statistical evidence for a linear dependence on $H$ within our framework. Therefore, to avoid a potentially misleading emphasis on an unsupported trend, we do not list the regression relations for these parameters. For completeness, we still provide the best-fit slope and intercept, as well as Pearson’s $r$, reduced $\chi^{2}$, and $\DeltaAIC/Δ\DeltaBIC$ in Table 2, enabling readers to directly compare the linear and constant models.
The correlation between $\log{P_{pk1}}$ and $H$ is
$\log{P_{pk1}}=(0.50\pm 0.04)\times H+(-0.02\pm 0.06),$ (12)
where $P_{pk1}$ is the peak photon flux in the time bin $64$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The Pearson’s $r$ is $0.48^{+0.004}_{-0.05}$, the reduced $\chi^{2}$ is $28.49$. The $\Delta AIC$ and $\Delta BIC$ are 41.76 and 38.67, respectively. The scatter plot is in Fig. 9.
The correlation between $\log{P_{pk2}}$ and $H$ is
$\log{P_{pk2}}=(0.61\pm 0.05)\times H+(-0.23\pm 0.07),$ (13)
where $P_{pk2}$ is the peak photon flux in the time bin $256$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The Pearson’s $r$ is $0.48^{+0.01}_{-0.03}$, the reduced $\chi^{2}$ is $51.92$. The $\Delta AIC$ and $\Delta BIC$ are 41.55 and 38.45, respectively. The scatter plot is in Fig. 10.
The correlation between $\log{P_{pk3}}$ and $H$ is
$\log{P_{pk3}}=(0.76\pm 0.07)\times H+(-0.49\pm 0.09),$ (14)
where $P_{pk3}$ is the peak photon flux in the time bin $1024$ ms of 10100010-1000 keV and is in units of ph cm-2s-1. The Pearson’s $r$ is $0.49^{+0.002}_{-0.03}$, the reduced $\chi^{2}$ is $72.54$. The $\Delta AIC$ and $\Delta BIC$ are 41.92 and 38.82, respectively. The scatter plot is in Fig. 11.
The correlation between $\log{T_{50}}$ and $H$ is
$\log{T_{50}}=(-0.93\pm 0.06)\times H+(2.18\pm 0.05),$ (15)
where $T_{50}$ is the burst duration containing $25\%-75\%$ of total counts and is in units of s. The Pearson’s $r$ is $0.64^{+0.03}_{-0.002}$, the reduced $\chi^{2}$ is $54.23$. The $\Delta AIC$ and $\Delta BIC$ are 82.94 and 79.85, respectively. The scatter plot is in Fig. 12.
The correlation between $\log{T_{90}}$ and $H$ is
$\log{T_{90}}=(-0.76\pm 0.05)\times H+(2.43\pm 0.04),$ (16)
where $T_{90}$ is the burst duration containing $5\%-95\%$ of total counts and is in units of s. The Pearson’s $r$ is $0.63^{+0.03}_{-0.004}$, the reduced $\chi^{2}$ is $41.82$. The $\Delta AIC$ and $\Delta BIC$ are 80.59 and 77.51, respectively. The scatter plot is in Fig. 13.
[…]
Parameters vs. $H$ Slope Intercept Pearson’s $r$ Reduced $\chi^{2}$ $\Delta AIC$ $\Delta BIC$
$\alpha$ $0.48\pm 0.05$ $1.83\pm 0.06$ $0.0033^{+0.03}_{-0.03}$ 37.41 -1.78 -4.88
$\beta$ $0.12\pm 0.03$ $2.23\pm 0.06$ $0.02^{+0.07}_{-0.09}$ 5.72 -337.55 -340.11
$\log{E_{peak}}$ $0.35\pm 0.04$ $1.83\pm 0.06$ $0.09^{+0.02}_{-0.04}$ 55.84 -0.69 -3.66
$\log{F_{g}}$ $1.43\pm 0.24$ $0.39\pm 0.27$ $0.13^{+0.01}_{-0.03}$ 117.76 0.74 -2.34
$\log{F_{pk}}$ $19\pm 0.12 0.0$$07\pm 0.18$ $0.24^{+0.07}_{-0.05}$ 29.13 -1.41 -1.97
HRHR $2.92\pm 0.58$ $04\pm 0.52 0.0$$01^{+0.02}_{-0.02}$ 99.90 -1.97 -5.05
$\log{P_{pk1}}$ $50\pm 0.04 0.0$$02\pm 0.06$ $0.48^{+0.004}_{-0.05}$ 28.49 41.76 38.67
$\log{P_{pk2}}$ $0.61\pm 0.05$ $0.23\pm 0.07$ $0.48^{+0.01}_{-0.03}$ 51.92 41.55 38.45
$\log{P_{pk3}}$ $0.76\pm 0.07$ $0.49\pm 0.09$ $0.49^{+0.002}_{-0.03}$ 72.54 41.92 38.82
$\log{T_{50}}$ $0.93\pm 0.06$ $2.18\pm 0.05$ $0.64^{+0.03}_{-0.002}$ 54.23 82.94 79.85
$\log{T_{90}}$ $0.76\pm 0.05$ $2.43\pm 0.04$ $0.63^{+0.03}_{-0.004}$ 41.82 80.59 77.51
$\rm{V}$ $01\pm 0.16 0.0$$04\pm 0.22$ 0.39-0.39
Table 2: Correlation coefficient results.
Figure 3: Scatter plot for $\alpha$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 4: Scatter plot for $\beta$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 5: Scatter plot for $\log{E_{peak}}$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 6: Scatter plot for $\log{F_{g}}$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 7: Scatter plot for $\log{F_{pk}}$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 8: Scatter plot for $HR$ and Hurst index. For this parameter, the AIC/BIC model comparison favors the null (constant) model over the linear model; therefore, no regression line is shown.
The description of each parameter is in Section 2.2.
Figure 9: Scatter plot for $\log{P_{pk1}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk1}}=(0.50\pm 0.04)\times H+(-0.02\pm 0.06)$. The description of each parameter is in Section 2.2.
Figure 10: Scatter plot for $\log{P_{pk2}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk2}}=(0.61\pm 0.05)\times H+(-0.23\pm 0.07)$. The description of each parameter is in Section 2.2.
Figure 11: Scatter plot for $\log{P_{pk3}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{P_{pk3}}=(0.76\pm 0.07)\times H+(-0.49\pm 0.09)$. The description of each parameter is in Section 2.2.
Figure 12: Scatter plot for $\log{T_{50}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{T_{50}}=(-0.93\pm 0.06)\times H+(2.18\pm 0.05)$. The description of each parameter is in Section 2.2.
Figure 13: Scatter plot for $\log{T_{90}}$ and Hurst index. The solid line is our fit result. The relation for the solid line is $\log{T_{90}}=(-0.76\pm 0.05)\times H+(2.43\pm 0.04)$. The description of each parameter is in Section 2.2.
Figure 14: Scatter plot for $V$ and Hurst index. The description of each parameter is in Section 2.2.
The correlation analysis between the Hurst index ($H$) and the 12 GRB observables is summarized in Table 2. The complete set of best-fit linear relations is given in Eqs. (12)–(16), and the corresponding scatter plots are shown in Figs. 314. In addition to the Pearson coefficient, we report the reduced $\chi^{2}$ from the linear fits and the information-criterion differences ($\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$) comparing the linear model against a null (no-trend) model (Table 2).
Overall, the strongest (anti-)correlations are found for the temporal parameters $T_{50}$ and $T_{90}$, with Pearson coefficients $r=-0.64^{+0.03}_{-0.002}$ and $r=-0.63^{+0.03}_{-0.004}$, respectively. Consistent with these coefficients, both $\log T_{50}$ and $\log T_{90}$ yield the largest positive $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ among all parameters, indicating that a linear trend is strongly preferred over the null model for these duration-related observables.
For the flux-related quantities, we find clear positive associations between $H$ and the peak photon flux proxies: $\log P_{pk1}$, $\log P_{pk2}$, and $\log P_{pk3}$ all show moderate positive correlations with $r\simeq 0.48$$0.49$ (Table 2). The corresponding $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ are also positive and similar for the three time bins, supporting the presence of a linear trend. By contrast, for both $\log F_{g}$ and $\log F_{pk}$ we find no evidence that a linear trend with $H$ is required by the data. The Pearson coefficients are $r=0.13^{+0.01}_{-0.03}$ for $\log F_{g}$ and $r=-0.24^{+0.07}_{-0.05}$ for $\log F_{pk}$, and the information-criterion comparison ($\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$) favors the null model over the linear model in both cases. Accordingly, we do not emphasize linear regression relations; for completeness, the linear-fit coefficients together with the corresponding $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ are listed in Table 2.
For the spectral- and hardness-related parameters, the data show no compelling evidence for a linear dependence on $H$. Consistent with the near-zero Pearson coefficients, information-criterion model comparison further indicates that the null model is preferred over the linear model for these parameters; accordingly, we do not display linear regression lines in the corresponding figures. In particular, the Pearson coefficient for $\alpha$ is consistent with zero, with $r=-0.0033^{+0.03}_{-0.03}$, and $HR$ is also consistent with no correlation, with $r=-0.01^{+0.02}_{-0.02}$. The variability index $V$ shows a negative Pearson coefficient r=0.39r=-0.39; however, given the absence of a well-defined $1\sigma$ measurement uncertainty for $V$ in the catalog, we do not apply our likelihood-based $\Delta\mathrm{AIC}/\Delta\mathrm{BIC}$ model comparison or Monte-Carlo uncertainty propagation for $r$ to this parameter in order to maintain statistical consistency.
Finally, we note that the reduced $\chi^{2}$ values for several relations are substantially larger than unity (Table 2), indicating that the observed scatter is not fully explained by the quoted measurement errors under a simple linear description. We therefore interpret the Pearson coefficient and $\Delta\mathrm{AIC}/\Delta\mathrm{BIC}$ as complementary diagnostics: $r$ characterizes the monotonic association strength, while the goodness-of-fit and information criteria quantify whether a linear model provides an adequate and parsimonious description of the data.
5 Conclusions and Discussion
In this work, we applied detrended fluctuation analysis (DFA) to the prompt-emission light curves of 163 long-duration BATSE GRBs and extracted the Hurst index $H$ as a quantitative descriptor of the scaling behavior of the fluctuation function. We then examined how $H$ relates to twelve commonly used GRB observables compiled by Wang et al. (2020), and summarized the correlation coefficients and linear-fit diagnostics in Table 2.
Our most robust empirical result is a strong anti-correlation between $H$ and the duration-related parameters: $\log T_{50}$ and $\log T_{90}$ yield Pearson coefficients of $r=-0.64^{+0.03}_{-0.002}$ and $r=-0.63^{+0.03}_{-0.004}$, respectively, and both relations are strongly preferred over a null (no-trend) model according to $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ (Table 2). This indicates that GRBs with longer durations tend to exhibit smaller DFA-derived $H$, i.e., weaker persistence in the scaling behavior of their prompt light curves. One plausible interpretation is that longer bursts may involve a larger number of emission episodes and/or a more heterogeneous superposition of variability components, which reduces the apparent long-range persistence captured by a single global DFA slope.
In the flux-related group, we find moderate positive correlations between $H$ and the peak photon flux proxies, with $r\simeq 0.48$$0.49$ for $\log P_{pk1}$, $\log P_{pk2}$, and $\log P_{pk3}$, and consistently positive $\Delta\mathrm{AIC}$ and $\Delta\mathrm{BIC}$ favoring a linear trend (Table 2). This suggests that bursts with more intense prompt emission (as traced by peak photon flux) tend to display larger $H$ values. Notably, within our sample, the correlation strengths for $P_{pk1}$$P_{pk3}$ are comparable rather than showing a clear monotonic weakening from 64 ms to 1024 ms, indicating that the association between $H$ and peak intensity is not confined to a single peak-flux timescale in the Wang et al. compilation.
[…]
Several caveats should be considered when interpreting these trends. First, many reduced $\chi^{2}$ values in Table 2 are substantially larger than unity, implying that a simple linear relation does not fully account for the scatter under the quoted measurement errors; intrinsic dispersion and/or unmodeled heterogeneity across the GRB population is likely important. Second, our DFA implementation adopts uniform binning and a single detrending order, and we fit a single slope over all adopted scales; GRBs with scale-dependent behavior (e.g., crossovers) may not be optimally characterized by a single global $H$. Third, all results are based on BATSE DISCSC light curves and the parameter definitions adopted by Wang et al. (2020); correlations may change with broader energy coverage, different temporal sampling, or alternative catalog constructions. Finally, for a small subset of bursts with truncated coverage ($t_{\rm end}^{\rm data}<T_{90}$) and/or collapsed scale range ($s_{\max}<s_{\min}$), the DFA fit is less constrained and thus expected to carry larger statistical uncertainties, reducing their leverage in sample-level correlations.
From Figs. 611, one can see a weak signature indicating that the sample can be classified into two classes in the Hurst index and flux plane. This phenomenon appears in all fluences, peak flux, and peak photon flux. It is similar to the Amati relation (Amati et al., 2002) and the Ghirlanda relation (Ghirlanda et al., 2004), which can be used to distinguish between the collapsar origin and the binary merging origin. However, as the samples are all long GRBs, it should not be the case for differentiating between long and short GRBs. This may provide another criterion for the classification of GRB sub-classes. These sub-classes might be used for those bursts with redshift measurements in future work. Then the physical quantities, such as luminosity and total energy, can be included, and the underlying origin of the sub-classes might be discovered.
Future work can extend this study by applying the same pipeline to broader-band datasets (e.g., Fermi/GBM), exploring time-resolved or scale-dependent DFA slopes within individual bursts, and comparing the observed HH–duration and HH–peak-flux trends with predictions from physically motivated prompt-emission simulations. In addition, we will quantify these potential limitations via robustness checks that repeat the correlation tests under stricter thresholds on the minimum time-series length and the minimum number of usable scales.
In summary, the DFA-derived Hurst index $H$ shows its clearest connections to prompt-emission timescales (durations) and intensity indicators (peak photon fluxes), while the standard spectral parameters provide no comparable evidence for a linear dependence on $H$ in the present sample. These empirical relations provide a quantitative link between the statistical structure of prompt light-curve variability and a subset of key GRB observables and motivate further multi-instrument and time-resolved investigations.
Acknowledgments
We thank the anonymous reviewer for his/her comments which enhanced this work enormously. We also thank Shuang-Xi Yi for useful comments and Cheng-Jie Sun for useful discussions. This work was supported by the National SKA Program of China (2022SKA0130100). The computation was performed on the HPC Platform of Huazhong University of Science and Technology.
Appendix A List of Hurst Indices
Table 3: List of Hurst Indices.
Trigger number GRB name Hurst index
107 GRB 910423 $0.42\pm 0.04$
109 GRB 910425A $0.40\pm 0.06$
110 GRB 910425B $0.22\pm 0.02$
111 GRB 910426 $0.14\pm 0.11$
114 GRB 910427 $0.50\pm 0.08$
121 GRB 910429 $0.42\pm 0.11$
130 GRB 910430 $0.27\pm 0.32$
133 GRB 910501 $0.46\pm 0.04$
143 GRB 910503 $2.08\pm 0.07$
148 GRB 910505 $0.17\pm 0.03$
160 GRB 910507 $1.84\pm 0.04$
171 GRB 910509 $1.10\pm 0.08$
204 GRB 910517B $0.31\pm 0.15$
211 GRB 910518B $0.22\pm 0.02$
214 GRB 910521B $2.41\pm 0.10$
219 GRB 910522 $1.57\pm 0.37$
222 GRB 910523 $0.23\pm 0.27$
223 GRB 910523B $0.69\pm 0.13$
226 GRB 910525 $0.26\pm 0.04$
228 GRB 910526B $1.62\pm 0.65$
235 GRB 910528 $0.79\pm 0.09$
237 GRB 910529 $0.06\pm 0.10$
249 GRB 910601 $3.05\pm 0.08$
257 GRB 910602 $0.24\pm 0.10$
288 GRB 910607B $0.46\pm 0.07$
332 GRB 910612 $1.28\pm 0.27$
351 GRB 910614B $0.42\pm 0.07$
394 GRB 910619 $0.44\pm 0.04$
398 GRB 910620 $1.44\pm 0.07$
404 GRB 910621B $0.19\pm 0.05$
408 GRB 910621 $0.26\pm 0.21$
414 GRB 910622 $1.24\pm 0.56$
451 GRB 910627 $2.65\pm 0.03$
465 GRB 910629B $0.75\pm 0.32$
467 GRB 910629 $2.35\pm 0.19$
469 GRB 910630 $2.15\pm 0.28$
472 GRB 910701 $0.42\pm 0.06$
473 GRB 910702B $0.90\pm 0.11$
501 GRB 910708 $1.41\pm 0.10$
503 GRB 910709 $1.74\pm 0.42$
516 GRB 910712 $1.50\pm 0.18$
540 GRB 910715 $1.05\pm 0.16$
548 GRB 910718B $1.86\pm 0.07$
559 GRB 910721B $1.03\pm 0.16$
563 GRB 910721 $1.78\pm 0.12$
577 GRB 910725B $0.18\pm 0.03$
591 GRB 910730B $0.33\pm 0.40$
594 GRB 910730 $1.75\pm 0.10$
606 GRB 910802B $1.54\pm 0.10$
630 GRB 910805 $0.70\pm 0.14$
647 GRB 910807 $1.37\pm 0.13$
658 GRB 910809C $1.33\pm 0.09$
659 GRB 910809B $0.32\pm 0.10$
660 GRB 910809 $2.33\pm 0.15$
673 GRB 910813 $0.77\pm 0.12$
676 GRB 910814C $0.85\pm 0.14$
678 GRB 910814A $0.32\pm 0.07$
680 GRB 910815 $1.57\pm 0.16$
685 GRB 910816B $0.77\pm 0.23$
686 GRB 910816 $1.38\pm 0.17$
692 GRB 910818 $0.22\pm 0.19$
704 GRB 910821 $0.81\pm 0.18$
717 GRB 910823 $0.93\pm 0.10$
741 GRB 910829 $0.85\pm 0.01$
753 GRB 910903 $0.77\pm 0.04$
761 GRB 910905 $0.55\pm 0.24$
764 GRB 910907 $1.10\pm 0.10$
773 GRB 910908 $1.07\pm 0.33$
795 GRB 910914 $1.19\pm 0.07$
815 GRB 910923 $0.94\pm 0.11$
816 GRB 910925 $0.46\pm 0.18$
820 GRB 910926 $1.15\pm 0.10$
824 GRB 910926B $0.90\pm 0.06$
825 GRB 910927B $0.10\pm 0.08$
829 GRB 910927 $2.32\pm 0.18$
840 GRB 910930B $0.61\pm 0.33$
841 GRB 910930 $1.99\pm 0.56$
869 GRB 911005 $0.38\pm 0.18$
907 GRB 911016 $0.37\pm 0.08$
927 GRB 911024 $1.23\pm 0.12$
938 GRB 911026 $1.34\pm 0.12$
946 GRB 911027B $0.56\pm 0.07$
973 GRB 911031A $0.81\pm 0.14$
1009 GRB 911106B $0.48\pm 0.03$
1036 GRB 911110 $0.30\pm 0.09$
1039 GRB 911111C $1.10\pm 0.13$
1042 GRB 911111 $1.53\pm 0.14$
1046 GRB 911111B $0.77\pm 0.05$
1085 GRB 911118A $3.03\pm 0.11$
1086 GRB 911118B $0.67\pm 0.14$
1087 GRB 911119B $1.03\pm 0.15$
1156 GRB 911209C $0.29\pm 0.05$
1157 GRB 911209 $0.01\pm 0.09$
1159 GRB 911210 $2.30\pm 0.82$
1167 GRB 911213 $0.84\pm 0.10$
1192 GRB 911217B $0.58\pm 0.10$
1196 GRB 911219B $0.26\pm 0.05$
1197 GRB 911219 $1.40\pm 0.09$
1200 GRB 911221 $1.34\pm 0.06$
1213 GRB 911224C $0.90\pm 0.09$
1218 GRB 911225 $1.65\pm 0.15$
1235 GRB 911227 $0.44\pm 0.09$
1244 GRB 911228 $1.05\pm 0.68$
1279 GRB 920105 $1.27\pm 0.09$
1288 GRB 920110A $0.23\pm 0.03$
1291 GRB 920110B $1.73\pm 0.01$
1384 GRB 920210C $0.46\pm 0.11$
1385 GRB 920210 $1.02\pm 0.16$
1390 GRB 920212 $0.69\pm 0.03$
1396 GRB 920214 $1.63\pm 0.13$
1406 GRB 920216 $1.68\pm 0.09$
1419 GRB 920218 $0.86\pm 0.27$
1425 GRB 920221 $2.68\pm 0.18$
1432 GRB 920224 $0.81\pm 0.10$
1440 GRB 920226 $2.64\pm 0.10$
1446 GRB 920227C $0.94\pm 0.07$
1447 GRB 920227B $1.68\pm 0.18$
1449 GRB 920228 $1.60\pm 0.11$
1452 GRB 920229B $1.24\pm 0.17$
1456 GRB 920301 $0.09\pm 0.06$
1458 GRB 920302B $0.06\pm 0.09$
1467 GRB 920307 $1.66\pm 0.15$
1468 GRB 920308A $0.50\pm 0.18$
1472 GRB 920310 $1.61\pm 0.20$
1574 GRB 920430 $0.11\pm 0.05$
1578 GRB 920502 $1.99\pm 0.08$
1579 GRB 920502C $1.22\pm 0.28$
1580 GRB 920503 $1.49\pm 0.15$
1586 GRB 920505 $0.59\pm 0.22$
1590 GRB 920509 $0.55\pm 0.11$
1601 GRB 920511A $0.33\pm 0.07$
1604 GRB 920511B $0.79\pm 0.08$
1606 GRB 920513 $0.31\pm 0.13$
1609 GRB 920517 $3.42\pm 0.15$
1611 GRB 920518 $1.09\pm 0.10$
1614 GRB 920520 $0.28\pm 0.06$
1623 GRB 920524 $0.08\pm 0.03$
1625 GRB 920525 $3.34\pm 0.11$
1626 GRB 920525C $1.78\pm 0.43$
1628 GRB 920526 $1.64\pm 0.08$
1642 GRB 920609 $0.13\pm 0.08$
1646 GRB 920613 $0.17\pm 0.07$
1652 GRB 920617 $0.76\pm 0.14$
1653 GRB 920617C $0.22\pm 0.39$
1655 GRB 920618 $0.03\pm 0.14$
1656 GRB 920619 $0.18\pm 0.07$
1657 GRB 920619B $0.99\pm 0.01$
1660 GRB 920620 $0.34\pm 0.17$
1661 GRB 920620C $0.79\pm 0.09$
1663 GRB 920622 $3.10\pm 0.08$
1667 GRB 920624 $0.29\pm 0.13$
1676 GRB 920627 $0.15\pm 0.17$
1687 GRB 920707 $0.55\pm 0.08$
1693 GRB 920710 $0.79\pm 0.17$
1922 GRB 920912 $2.00\pm 0.07$
1924 GRB 920913 $1.27\pm 0.11$
1956 GRB 920925 $1.93\pm 0.16$
1967 GRB 921001 $1.55\pm 0.08$
1982 GRB 921008 $0.73\pm 0.07$
1989 GRB 921015 $0.58\pm 0.04$
1991 GRB 921017 $0.54\pm 0.07$
1993 GRB 921021 $2.04\pm 0.22$
1997 GRB 921022 $1.48\pm 0.53$